OpenStudy (odie8136):

Write an equation of a line whose graph is parallel to the graph of y = 3x – 10

OpenStudy (3mar):

May I help?

OpenStudy (odie8136):

yes you may 3mar

OpenStudy (3mar):

The form of line equation requires two inputs: - the slope of the line (which could be gotten from the parallel line) - a point on the line itself or the intersect position with y-axis.

OpenStudy (3mar):

How many input did you get?

OpenStudy (odie8136):

i don't understand

OpenStudy (3mar):

you need the slope of the line + one point it passes through \[\frac{ y-y_1 }{ x-x_1 }=m\]

OpenStudy (3mar):

or you need two points on the line: \[\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{ y-y_1 }{ x-x_1 }\]

OpenStudy (3mar):

or you need the slope + intersect with y-axis \[y=mx+b\] m: slope b: intersection between y-axis and the line

OpenStudy (3mar):

The slope is easily calculated by the parallel line. "Any two parallel lines have the same slope." so the slope of the two lines is \[m=3\]

OpenStudy (3mar):

We need now one more given in order to write the line equation.

OpenStudy (3mar):

Are you with me?